It takes 20 minutes to set up a candy making machine. Once the machine is set up, it produces 25 candies per minute. Use an inequality to find the number of candies that can be produced in 2 hours if the machine has not been set up.
Answer by Theo(13342)

2 hours * 60 minutes per hour is equal to 120 minutes.

it takes 20 minutes to set up the machine, which leaves 100 minutes to produce candies.

the machine can produce 25 candies per minute * 100 minutes = 2500 candies.

your formula can be set up as follows:

c = 25 * (m - s)

c is the number of candies to produce.
m is the number of minutes you have available.
s is the time required to set up the machine in minutes.

if the machine is already set up, then s = 0 and the formula becomes:

c = 25 * m.

if the machine is not already set up, then s = 20 and the formula becomes:

c = 25 * (m - 20)

it's the second formula that we used.

m was equal to 2 hours * 60 minutes per hour = 120 minutes.

the formula became c= 25 * (120 - 20) which became c = 25 * 100 which became c = 2500.